Helppp
Find the angles between the lines AB and AC as shown below. B(-3,5,7) C(5,7, 4) A(4.31
Answer #1

For this problem, we can represent the lines AB and AC as vectors and further the cosine formula for vector multiplication can be used to get the value of the required angle.

The steps to be followed to obtain the angle are as follows:-

We know, AB AC = LABIAC I cos o © - sos AB AC - 1 BIJAC) -0 here, AB = (-3-4) + (5-3) + (7-2) h = -7ût 29 + 5 TAB) = (-7)+(2

Thus the required angle between lines AB and AC = 74.230